OpenStudy (anonymous):

HELP ME PLEASE !#$#!@#!@#@!$%$~! Part A: Divide (4x4y3 + 8x3y2 - 12x2y - 16x2y4) by -4x2y. Show your work, and justify each step. (6 points) Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? (2 points) Part C: What is the degree and classification of the polynomial you got in Part A?

OpenStudy (anonymous):

I thought I helped.

OpenStudy (anonymous):

NEED PART B AND C

OpenStudy (mertsj):

B All the x exponents in the numerator would be 1 greater

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

can you please help with part c

OpenStudy (mertsj):

I don't know what you have for part A or how can I?

OpenStudy (anonymous):

was part a anser -56xy2-8xy... lol i dont know really suck at math

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

http://openstudy.com/users/amoodarya

OpenStudy (amoodarya):

(4x4y3 + 8x3y2 - 12x2y - 16x2y4) 4((x4y3 + 2x3y2 - 3x2y - 4x2y4)= 4 x^2 (x^2y^3 +2xy^2 -3y-4y^4)= 4 x^2 y (x^2y^2 +2xy -3-4y^3)= now divide by -4x^2y 4 x^2 y (x^2y^2 +2xy -3-4y^3) --------------------------- = - (x^2y^2 +2xy -3-4y^3) -4x^2 y

OpenStudy (anonymous):

thanks!!!!!!!!!!!

OpenStudy (anonymous):

also one last thing can you please help with part c

OpenStudy (amoodarya):

degree of (4x4y3 + 8x3y2 - 12x2y - 16x2y4) for x is 4 because of x^4 for y is 4 because of y^4

OpenStudy (anonymous):

thanks so much